# Register Always Matters: Analysis of LLM Pretraining Data Through the Lens of Language Variation

### v1

@@ 5.1 Average performance
The results for individual register models, averaged over all benchmarks, are presented in Figure 1 and detailed in Appendix B Figure 4 and Appendix C. The evaluation shows striking differences in accuracy between the models, demonstrating that the register of the pretraining data has a substantial effect on model performance. We find that the model trained on the full HPLT v2 data outperforms ones trained on individual registers, demonstrating that a mix of registers is important for LLM training. The performance ranking of individual registers generally aligns with intuitions about their content and what is typically considered high-quality pretraining data: registers like Lyrical (LY), Machine Translation (MT), and Informational Persuasion (IP) yield worse-performing models, while models trained on registers containing informational content and instructions, such as Description (dtp) and How-to-Instructions (HI), perform the best. Surprisingly, Opinion (OP) yields the 4th best performing model. Both Informational Description (IN) and Narrative (NA) models perform worse than How-to-Instructions (HI), Description (dtp) and Opinion (OP). Following these, with a noticeable performance drop, are the models trained on Interactive Discussion (ID), Spoken (SP) and remarkably, the subregister News (ne), consisting of news articles, which performs among the worst, in contrast with work by de la Rosa et al. (2025).
@@
As opposed to the main register models, the hybrid class model Instructive-Informational (HI-IN) outperforms the model trained on the full HPLT v2 data. This shows that hybrid documents do not simply result in averaged performance between the two hybridized registers, but can show better performance than either alone. This finding motivated our second experiment investigating register combinations. The subregisters News (ne) (subregister of Narrative) and Description (dtp) (subregister of Informational Description) show divergent patterns when compared to their parent registers: News drastically underperforms its main-level register Narrative, while Description achieves higher accuracy than Informational Description. We analyze differences between main-subregister pairs in more detail in the following section.
Figure 1: Average normalized accuracy over all benchmarks for individual register models and HPLT v2 model, shown as a range from first checkpoint (2 billion tokens) to final checkpoint (100 billion tokens), in order of final checkpoint performance. The first checkpoint presented as gray, last checkpoint with color.
Figure 2: Averaged normalized accuracy over all benchmarks for combination models, HPLT v2 and FineWeb models. Change visualized as range from first checkpoint (2 billion tokens) to final checkpoint (100 billion tokens), with the first checkpoint presented as gray, last checkpoint with color.
@@
Figure 3: Change of accuracy from first to final checkpoint on individual benchmarks shown as a range, with gray indicating the first checkpoint and colors indicating last checkpoint. The random-guess threshold is shown as a gray line in cases where at least one model falls below it. Bars and legend shown in order of average accuracy.
@@
As the results show, pretraining a model with data from only one register class does not provide improvements compared to training with the full dataset. This aligns with the established notion that pretraining data should have variability to yield capable models. The results for our combination model experiment are shown in Figure 2, and again with more detail in Appendix 5. The results clearly support the above idea: all combination models outperform the individual register models and the baseline HPLT v2 model. Additionally, the best models in our study (HI-IN-HI-dtp-OP and HI-IN-HI-dtp-OP-NA) reach higher average accuracy than the model trained on the FineWeb dataset. HI-IN-HI-dtp model shows much worse performance than HI-IN-HI-dtp-OP, which further includes Opinion (OP). This suggests that opinionated texts provide valuable pretraining data and are key to high model performance, at least as measured by the selected benchmarks. Further adding the Narrative (NA) dataset, as seen in the HI-IN-HI-dtp-OP-NA model, slightly worsens performance, though the difference is small. Significant drops occur when including Interactive Discussion (ID) and Spoken (SP) in models HI-IN-HI-dtp-OP-NA-ID and HI-IN-HI-dtp-OP-NA-ID-SP, indicating either that these registers contain qualities that harm the abilities measured on these benchmarks, or that they do not provide benefits that outweigh the reduced proportion of Instructive-Informational (HI-IN), Description (dtp), and Opinion (OP) in the training mix.
@@ 5.2 Per task
The results for each benchmark separately are shown in Figure 3 and detailed further in Appendix B Figure 6. The differences between benchmarks are remarkable, both in the performance ranking of models and in the way accuracy trends with respect to the number of training tokens. This shows that registers have strengths in different tasks and that the selection of benchmarks has a tangible impact on the average results.
@@
Most benchmarks display the monotonicity condition that influenced the selection of Penedo et al. (2024a), as they tend towards increased accuracy with more training tokens. A clear outlier is the SIQA benchmark, where increases in accuracy are very small, with the largest increase seen in Narrative (NA) and Instructive-Informational (HI-IN) models, both increasing approximately by 0.03 during the training. This is further seen in Appendix B Figure 6.
@@
Generally, the trends observed in Section 5.1 hold. The models with the lowest average accuracy—Lyrical (LY), Machine Translation (MT) and Informational Persuasion (IP)—perform poorly on all benchmarks, and, notably, worse than random guess on ARC Challenge. The model for subregister News (ne), despite weak average performance, improves on the MMLU and WinoGrande benchmarks, while Spoken (SP) enhances results on CommonsenseQA. The differences between benchmarks become more pronounced with models trained on Narrative (NA), Informational Description (IN), Opinion (OP), How-to-Instructions (HI), and Interactive Discussion (ID), which display great variation in ranking across tasks, highlighting their strengths as pretraining material. The Narrative model performs well on SIQA and WinoGrande but barely surpasses the random baseline on ARC Challenge. Informational Description trumps other models on MMLU and both ARC benchmarks but ranks low to mediocre elsewhere. The model trained on the How-to-Instructions dataset outperforms almost all other models on PIQA and HellaSwag. This might be because PIQA measures reasoning in physical tasks, which aligns with the instructions related to physics included in How-to-Instructions. On the other hand, How-to-Instructions ranks among the worst on MMLU and CommonsenseQA. The Opinion model excels in WinoGrande, SIQA and CommonsenseQA. This signal also appears in the combination models’ performance on CommonsenseQA: including Opinion in the training data greatly increases accuracy on this benchmark, unlike the HI-IN-HI-dtp model, which excludes Opinion.
@@
The hybrid Instructive-Informational (HI-IN) model generally performs between individual How-to-Instructions (HI) and Informational Description (IN) models, and is notably closer in performance to whichever yields better accuracy. Exceptions to this pattern are OpenBookQA, CommonsenseQa, and WinoGrande, where the Instructive-Informational model shows better accuracy than either How-to-Instructions or Informational Description separately. Specifically in the case of WinoGrande, this is surprising, as both How-to-Instructions and Informational Description struggle with this benchmark. From this result, we conclude that hybrids may contain qualities not found in individual registers and require further study. The News (ne) model trails behind Narrative (NA), except on the MMLU and ARC Challenge benchmarks. This is likely due to news articles covering topics related to MMLU tasks, such as politics, business, science, and environment. Similarly, the other main-subregister pair, Description (dtp) and Informational Description (IN), shows differences in performance across benchmarks, with the Description-model outshining Informational Description in ARC Easy, HellaSwag, PIQA, WinoGrande, and OpenbookQA. Informational Description, on the other hand, excels in ARC Challenge and MMLU. Specifically for the WinoGrande benchmark, the performance difference might be explained by the Description register containing more texts focused on things and people, thus resulting in better pronoun-resolution abilities.
@@
The combination models lead in performance on most benchmarks. On MMLU specifically, we observe a huge drop between the HI-IN-HI-dtp-OP model and others: including OP in the pretraining data increases performance, while removing OP or further adding NA greatly disrupts model performance on this benchmark. Notably, the models HI-IN-HI-dtp-OP and HI-IN-HI-dtp-OP-NA are trained with registers that individually perform well on different benchmarks: HI-IN on ARC Challenge, HI on Hellaswag and PIQA, dtp on ARC Easy, OpenBookQA and MMLU,OP on CommonsenseQA, and NA on SIQA and WinoGrande.
@@ 6 Conclusion
In this paper, we presented the first study investigating the effect of linguistic register on the downstream performance of small generative LLMs. Our findings show that register is an important explainer of LLM performance, and we were able to show surprising relationships between the register of pretraining data and model accuracy on different benchmarks. While using the whole dataset yielded better performance than limiting the pretraining data to any individual register class, we found specific combinations of registers that outperform the uncurated dataset and the FineWeb dataset. These findings can be used to increase pretraining data transparency and to optimize data selection methods, potentially addressing specific performance gaps by incorporating appropriate registers.
@@
Our experiments are conducted in English only. While register classification has shown multilingual prowess, our results might not generalize to other languages. We leave multilingual analysis for future work. The benchmarks used, though standard in studies of this type, do not fully measure all model capabilities, as discussed regarding the Lyrical class. Evaluating with more diverse benchmarks would provide more comprehensive results and might reveal strengths and drawbacks we were unable to uncover.
Bias from data curation methods is known to propagate to trained models (Rae et al., 2022; Gururangan et al., 2022), and register classification is likely not exempt from this issue. To mitigate bias, we used the deduplicated version of the HPLT v2 data. Although this version has been processed with Trafilatura, Penedo et al. (2024a) note that Trafilatura can leave “undesirable” content in the dataset, potentially affecting some registers more than others. However, our manual analysis of the datasets does not support this conclusion. The possibility of benchmark contamination also exists; register classification could systematically assign leaked benchmark material to specific classes, potentially explaining some performance differences.
@@ Acknowledgments
This project has received funding from the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC under decision number VN/3137/2024-OKM-6, the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101070350, and the Digital Europe Programme under grant agreement No. 101195233. The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the European Union. We wish to acknowledge CSC - IT Center for Science, Finland for computational resources.
@@
Training large language models is resource-demanding. This fact guided our decisions: we used small model sizes and trained only up to the point where clear conclusions can be drawn. In total, we used approximately 11,000 GPU hours per model. See Appendix A for calculation of the total computational cost. To further mitigate the environmental impact of training, we used LUMI supercomputer, which is among the most environmentally friendly supercomputers in the world [^3] https://lumi-supercomputer.eu/sustainable-future/. The models trained for this study are planned to be used in future research.
Appendix A Model training setup
[…]
@@ Appendix A Model training setup
Training each model took approximately 84 hours, amounting to 10,752 GPU hours per model with the node setup described above. Evaluation was carried out on a single GPU, and with 19 models, each with 50 checkpoints requiring approximately 20 minutes to evaluate, this added 300 GPU hours to the total computational cost. Other processing tasks, such as sampling and tokenization, required no GPU resources and used a comparatively negligible amount of CPU hours.
Appendix B Model performance by checkpoints
[…]
Appendix C Numerical results

### v2

@@ 5.1 Average performance
Figure 1: Individual register and HPLT v2 (deduplicated) model accuracies as a function of the number of training tokens. Rolling average over 6 billion tokens (3 adjacent checkpoints) applied for ease of reading overlapping lines, see Table 2 for numerical results for the final checkpoint. Dotted lines indicate training continuing over one epoch, legend in order of last checkpoint performance.
Figure 2: Combination model and baseline accuracies as a function of the number of training tokens. Rolling average over 6 billion tokens (3 adjacent checkpoints) applied for ease of reading overlapping lines, see Table 2 for numerical results for the final checkpoint. Legend in order of last checkpoint performance.
@@
The results for individual register models, averaged over all benchmarks, are presented in Figure 1 and Table 2, and detailed in Appendix D. The evaluation shows striking differences in accuracy between the models, demonstrating that the register of the pretraining data has a substantial effect on model performance. We find that the model trained on the full HPLT v2 data outperforms models trained on individual registers, demonstrating that a mix of registers is important for LLM training. The performance ranking of individual registers generally aligns with intuitions about their content and what is typically considered high-quality pretraining data: registers like Lyrical (LY), Machine Translation (MT), and Informational Persuasion (IP) yield worse-performing models, while models trained on registers containing informational content and instructions, such as Description (dtp) and How-to-Instructions (HI), perform the best. Surprisingly, Opinion (OP) yields the 4th best performing model. Both Informational Description (IN) and Narrative (NA) models perform worse than How-to-Instructions, Description and Opinion. Following these, with a noticeable performance drop, are the models trained on Interactive Discussion (ID), Spoken (SP) and remarkably, the subregister News (ne), consisting of news articles, which performs among the worst.
@@
As opposed to the main register models, the hybrid class model Instructive-Informational (HI-IN) outperforms the model trained on the full HPLT v2 data. This shows that hybrid documents do not simply result in averaged performance between the two hybridised registers, but can show better performance than either alone. This finding motivated our second experiment, investigating register combinations. The subregisters News (ne) (subregister of Narrative) and Description (dtp) (subregister of Informational Description) show divergent patterns when compared to their parent registers: News drastically underperforms its main-level register Narrative, while Description achieves higher accuracy than Informational Description. We analyse differences between main-subregister pairs in more detail in the following section.
Pretraining dataset Accuracy Pretraining dataset Accuracy
HI-IN-HI-dtp-OP 0.475 Opinion (OP) 0.447
HI-IN-HI-dtp-OP-NA 0.472 Narrative (NA) 0.441
HI-IN-HI-dtp 0.466 Informational Description (IN) 0.437
FineWeb 0.466 Interactive Discussion (ID) 0.431
HI-IN-HI-dtp-OP-NA-ID-SP 0.465 Spoken (SP) 0.422
Instructive-Informational (HI-IN) 0.465 News (ne) 0.418
HI-IN-HI-dtp-OP-NA-ID 0.464 Informational Persuasion (IP) 0.393
HPLT v2 (deduplicated, all registers) 0.457 Machine Translation (MT) 0.374
Description (dtp) 0.452 Lyrical (LY) 0.358
How-to-Instructions (HI) 0.447
Table 2: Results by the last checkpoint performance in numerical format. Note that the values differ slightly from the Figures 1 and 2 due averaging over adjacent checkpoints in the figures.
@@
Figure 3: Change of accuracy from first to final checkpoint on individual benchmarks shown as a range, with grey indicating the first checkpoint and colours indicating the last checkpoint. The random-guess threshold is shown as a grey vertical line in cases where at least one model falls below it. Bars and legend shown in order of average accuracy.
@@
As the results show, pretraining a model with data from only one register class does not provide improvements compared to training with the full dataset. This aligns with the established notion that pretraining data should have variability to yield capable models. The results for our combination model experiment are shown in Figure 2, Table 2, and again with more detail in Appendix D. The results clearly support the above idea: all combination models outperform the individual register models and the baseline HPLT v2 model. Additionally, the best models in our study (HI-IN-HI-dtp-OP and HI-IN-HI-dtp-OP-NA) reach higher average accuracy than the models trained on the HPLT v2 and FineWeb datasets. HI-IN-HI-dtp model shows much worse performance than HI-IN-HI-dtp-OP, which further includes Opinion (OP). This suggests that opinionated texts provide valuable pretraining data and are key to high model performance, at least as measured by the selected benchmarks. Further adding the Narrative (NA) dataset, as seen in the HI-IN-HI-dtp-OP-NA model, slightly worsens performance, though the difference is small. Significant drops occur when including Interactive Discussion (ID) and Spoken (SP) in models HI-IN-HI-dtp-OP-NA-ID and HI-IN-HI-dtp-OP-NA-ID-SP, indicating either that these registers contain qualities that harm the abilities measured on these benchmarks, or that they do not provide benefits that outweigh the reduced proportion of Instructive-Informational (HI-IN), Description (dtp), and Opinion (OP) in the training mix.
@@ 5.2 Per task
The results for each benchmark separately are shown in Figure 3 and detailed further in Appendix E. The differences between benchmarks are remarkable, both in the performance ranking of models and in the way accuracy trends with respect to the number of training tokens. This shows that registers have strengths in different tasks. Most benchmarks display the monotonicity condition that influenced the selection of Penedo et al. (2024a), as they tend towards increased accuracy with more training tokens. A clear outlier is the SIQA benchmark, where increases in accuracy are very small, with the largest increase seen in Narrative (NA) and Instructive-Informational (HI-IN) models, both increasing approximately by 0.03 during the training.
@@
Generally, the trends observed in Section 5.1 hold. The models with the lowest average accuracy—Lyrical (LY), Machine Translation (MT) and Informational Persuasion (IP)—perform poorly on all benchmarks, and, notably, worse than a random guess on ARC Challenge. The model for subregister News (ne), despite weak average performance, improves on the MMLU and WinoGrande benchmarks, while Spoken (SP) enhances results on CommonsenseQA. The differences between benchmarks become more pronounced with models trained on Narrative (NA), Informational Description (IN), Opinion (OP), How-to-Instructions (HI), and Interactive Discussion (ID), which display great variation in ranking across tasks, highlighting their strengths. The Narrative model performs well on SIQA and WinoGrande but barely surpasses the random baseline on ARC Challenge. Informational Description trumps other models on MMLU and both ARC benchmarks, but ranks low to mediocre elsewhere. The model trained on the How-to-Instructions dataset outperforms almost all other models on PIQA and HellaSwag. This might be because PIQA measures reasoning in physical tasks, which aligns with the instructions related to physics included in How-to-Instructions. On the other hand, How-to-Instructions ranks among the worst on MMLU and CommonsenseQA. The Opinion model excels in WinoGrande, SIQA and CommonsenseQA. This signal also appears in the combination models’ performance on CommonsenseQA: including Opinion in the training data greatly increases accuracy on this benchmark, unlike the HI-IN-HI-dtp model, which excludes Opinion.
@@
The hybrid Instructive-Informational (HI-IN) model generally performs between individual How-to-Instructions (HI) and Informational Description (IN) models, and is notably closer in performance to whichever yields better accuracy. Exceptions to this pattern are OpenBookQA, CommonsenseQA, and WinoGrande, where the Instructive-Informational model shows better accuracy than either How-to-Instructions or Informational Description separately. Specifically in the case of WinoGrande, this is surprising, as both How-to-Instructions and Informational Description struggle with this benchmark. From this result, we conclude that hybrids may contain qualities not found in individual registers and require further study. The News (ne) model trails behind Narrative (NA), except on the MMLU and ARC Challenge benchmarks. This is likely due to news articles covering topics related to MMLU tasks, such as politics, business, science, and the environment. Similarly, the other main-subregister pair, Description (dtp) and Informational Description (IN), shows differences in performance across benchmarks, with the Description-model outshining Informational Description in ARC Easy, HellaSwag, PIQA, WinoGrande, and OpenbookQA. Informational Description, on the other hand, excels in ARC Challenge and MMLU. Specifically for the WinoGrande benchmark, the performance difference might be explained by the Description register containing more texts focused on things and people, thus resulting in better pronoun-resolution abilities.
@@
The combination models lead in performance on most benchmarks. On MMLU specifically, we observe a huge drop between the HI-IN-HI-dtp-OP model and others: including Opinion in the pretraining data increases performance, while removing Opinion or further adding Narrative greatly disrupts model performance on this benchmark. Notably, the models HI-IN-HI-dtp-OP and HI-IN-HI-dtp-OP-NA are trained with registers that individually perform well on a wide range of benchmarks: Instructive-Informational on HellaSwag and ARC Easy, How-to-Instructions on Hellaswag and PIQA, Description on ARC Easy, ARC Challenge, and OpenBookQA, Opinion on CommonsenseQA, and Narrative on SIQA and WinoGrande.
@@ 6 Conclusion
In this paper, we presented the first study investigating the effect of linguistic register on the performance of small generative LLMs. Our findings show that register is an important explainer of LLM performance, and we were able to show surprising relationships between the register of pretraining data and model accuracy on different benchmarks. While using the whole dataset yielded better performance than limiting the pretraining data to any individual register class, we found specific combinations of registers that outperform the full uncurated dataset. These findings can be used to increase pretraining data transparency and to optimise data selection methods, potentially addressing specific performance gaps by incorporating appropriate registers.
@@
As a first study in the intersection of two broad lines of research – the linguistic registers and LLM training – our work leaves a number of questions incompletely answered. Our experiments are conducted in English only. While register classification has shown multilingual prowess, our results might not generalise to other languages, and we leave multilingual analysis for future work. The benchmarks used, though standard in studies of this type, do not fully measure all model capabilities, as discussed regarding the Lyrical class. Evaluating with more diverse benchmarks would provide more comprehensive results and might reveal strengths and drawbacks we were unable to uncover.
Bias from data curation methods is known to propagate to trained models (Rae et al., 2022; Gururangan et al., 2022), and register classification is not exempt from this issue. To mitigate bias, we used the deduplicated version of the HPLT v2 data. Although this version has been processed with Trafilatura, Penedo et al. (2024a) note that Trafilatura can leave “undesirable” content in the dataset, potentially affecting some registers more than others. The possibility of benchmark contamination also exists; register classification could systematically assign leaked benchmark material to specific classes, potentially explaining some performance differences.
@@ Acknowledgments
This project has received funding from the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC under decision number VN/3137/2024-OKM-6, the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101070350, the Digital Europe Programme under grant agreement No. 101195233, and the Research Council of Finland under grant No. 362459. The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the European Union. We wish to acknowledge CSC - IT Center for Science, Finland, for computational resources.
@@
Training large language models is resource-demanding. This fact guided our decisions: we used small model sizes and trained only up to the point where clear conclusions can be drawn. In total, we used approximately 11,000 GPU hours per model. See Appendix B for calculation of the total computational cost. To further mitigate the environmental impact of training, we used LUMI supercomputer, which is among the most environmentally friendly supercomputers in the world [^3] https://lumi-supercomputer.eu/sustainable-future/. The models trained for this study are planned to be used in future research.
Releasing our models publicly for reproducibility opens up the possibility of our models being used in contexts other than research. Although we do not evaluate the possible toxicity of the generations they produce, the model sizes are small, and they are trained only up to 100 billion tokens, which makes their generative capabilities weak and thus likely unsuitable for applications outside the scope of this study.
Reproducibility statement
We have prioritised reproducibility in our work and make all artefacts created as part of this study, including code, model training configurations, and the trained model checkpoints, available on https://huggingface.co/TurkuNLP/register-ablations under open licenses. Our work only makes use of data openly available from https://hplt-project.org/datasets/v2.0, and our model training setup is detailed in Appendix B.
Appendix A Examples of texts in register classes
Below we show examples of texts classified as selected registers. These texts were chosen randomly but with a focus towards brevity (with some texts additionally truncated), general topics, and non-toxic language use. We also omit new-line characters.
How-to-Instructions (HI)
Home & DIY How To Cut Your Own Hair At Home With countless non-essential businesses closing due to the government-implemented lockdowns, many people have been forced to be creative and resourceful in order to get by. Case in point, not everyone of us are lucky enough to be quarantined at home with a professional hairdresser. So we do the best we can – which often means cutting our own hair. […]
Interactive Discussion (ID)
I tried last night to compile a standalone which I’ve compiled approximately 1, 000, 000 times before, and have been met with a long series of very lengthy crash reports. As far as I can tell, something in the compiler is trying to install every framework on my computer, and I sometimes wind up with 1 gig+ .mxf files. Naturally, little dinky tests compile fine, and I can compile all of the modules that make up my patch, but not the whole thing at once. I did a super- clean Max reinstall, no third-party objects aside a couple of Lobjects which are up to date, and I’m stumped. The only difference I can think of since I compiled this a few weeks ago is an upgrade to 4.5. Anyone have any light to throw on this? Im happy to send crashlogs, but they’re big. 10.5.2, Max 4.6, Jitter 1.6.3 thanks, M
Machine Translation (MT)
Incredible Interior Design Sketches Good Looking Interior Designer Remarkable Interior Design Exceptional Interior Design Sketches Cozy Interior Design Sketches Chic Interior Design Sketches Cool Interior Design Sketches
News (ne)
Wal-Mart has signed a storm water settlement agreement with the Connecticut Department of Environmental Protection concerning alleged violations at 20 Wal-Mart stores and two SAM’S CLUB locations in the state. Under the agreement, Wal-Mart will pay \$600,000 in civil penalties for violations alleged to have taken place between 1996 and 2003. Wal-Mart also will contribute \$550,000 to two different supplemental environmental projects–\$500,000 to assist municipalities in addressing storm water issues, and \$50,000 for environmental projects in the Connecticut River Watershed. […]
Opinion (OP)
Once again a very nice stop for the night. Staff has been very well trained, all are friendly and helpful with one exception/the young man who checked us in. Breakfast was so appreciated after 12 nights at a Holiday Inn Express which was ok but does not have your choices. Lynne does an especially good job/appreciated all her help. They… More
Appendix B Model training setup
Models were trained following the setup of Penedo et al. (2024a). We used the GPT-2 tokeniser, Llama architecture and the same training settings: 1.71B parameters (1.82B with embeddings), sequence length of 2048 tokens, and a global batch size of $\sim$2 million tokens. Models were trained for 50,000 iterations, which amounts to 100 billion tokens. Training was done on LUMI supercomputer on 16 nodes, each with a single 64-core CPU and 4 MI250x GPUs with dual-GCD (graphics compute die). We used NVIDIA’s Megatron-LM  [^4] https://github.com/NVIDIA/Megatron-LM training framework instead of HuggingFace’s nanotron [^5] https://github.com/huggingface/nanotron framework used by Penedo et al. (2024a).
@@ Appendix B Model training setup
Training each model took approximately 84 hours, amounting to 10,752 GPU hours per model with the node setup described above, with average performance of 25 teraflops per second. Evaluation was carried out on a single GPU, and with 19 models, each with 50 checkpoints requiring approximately 20 minutes to evaluate, this added 300 GPU hours to the total computational cost. Other processing tasks, such as sampling and tokenisation, required no GPU resources and used a comparatively negligible amount of CPU hours.
Appendix C Comparison to other data curation schemes
As discussed in Section 2, many different training data selection schemes and tools have been introduced in the literature. To quantify the connection and possible similarities between our register scheme and three other publicly available quality classification tools, FineWeb-edu classifier (Penedo et al., 2024a), the DCLM quality classifier (Li et al., 2024), and NVIDIA’s NemoCurator Quality Classifier DeBERTa (He et al., 2023), we classify a sample of over 50 000 documents from our pretraining material with all four classifiers. We compare the selected tool to our register classification scheme by conducting a $\chi^{2}$-test to evaluate dependence and also report metrics pertaining to the possible overlap.
The FineWeb-edu classifier assigns documents an ordinal label ranging from 0 to 5, the NemoCurator a label 0, 1, or 2, and the DCLM classifier uses a binary label (0,1). For registers, we use the classes in Table 1 under “Individual registers”. Obtained results from a $\chi^{2}$-test (considering FineWeb-edu and NemoCurator’s labels nominal) show that all quality classification tools are dependent on the register classification results with $p<0.01$. This dependence between the classification schemes is to be expected, as all schemes are used as a proxy for “quality” in LLM pretraining. To evaluate the strength of this connection, we calculated Cramer’s V, which yielded the following results:
  • Registers – DCLM classifier: 0.136
  • Registers – FineWeb-edu classifier: 0.221
  • Registers – NemoCurator classifier: 0.419
In the context of machine learning, these values indicate a low to moderate connection between the classifiers, with the NemoCurator showing the greatest association, which can be considered somewhat substantial. To account for the two ordinal classification schemes, FineWeb-edu and NemoCurator, we also calculated the $\eta^{2}$-metric, which measures the amount of impact a variable has on the variance of another. These results were 0.173 and 0.273, for FineWeb-edu and NemoCurator, respectively, which similarly show a strong connection between the schemes but do not imply direct overlap.
Although the above analysis shows associations between the studied schemes, the register classes coincide with all classes of the three tools, with the exception of FineWeb-edu’s highest quality class 5, which only occurs in conjunction with register classes Instructive-Informational and Informational Description. We visualise this in Figures 4, 5, and 6. This means that our register scheme can be seen as mutually reinforcing when combined with other established quality labelling schemes. We specifically see from our results that the 3 tools we compared to the register scheme often assign high-quality labels to documents from registers Informational Description, Description, and the case of NemoCurator, News, but neglect classes that we found beneficial in this study, such as How-to-Instructions and Opinion. This highlights the value of using the non-binary but linguistically motivated and versatile scheme that the registers offer for curating data.
Figure 4: Visualised cross-classification for our register scheme and the DCLM quality classifier. DCLM label 0 corresponds to low quality and label 1 to high quality.
Figure 5: Visualised cross-classification for our register scheme and the FineWeb-edu classifier. FineWeb-edu label 0 corresponds to the lowest quality and label 5 to the highest quality.
Figure 6: Visualised cross-classification for our register scheme and the NVIDIA NemoCurator classifier. NemoCurator label 0 corresponds to the highest quality and label 2 to the lowest quality.
Appendix D Numerical results